Aluminum nitrite and ammonium chloride react to form aluminum chloride, nitrogen, and water. What mass of AlCl3 is present after 99.75 g aluminum nitrite and 84.94 g ammonium chloride react completely?
1
Expert's answer
2020-06-05T05:12:16-0400

Al(NO)3+3NH4Cl=AlCl3+3N2+6H20


"n(Al(NO)3)=\frac{m}{M}=\frac{99.75}{164.9965}=0.60"


"n(NH4Cl)=\frac{m}{M}=\frac{84.94}{53.4920}=1.59"


aluminum nitrite is in short supply, then we calculate it

n(AlCl3)=0.60


"m=n\times M=0.6\times 133.3405=80.00" g